# Layer-adaptive Expert Pruning for Pre-Training of Mixture-of-Experts Large Language Models

### v2

@@ Abstract
Although Mixture-of-Experts (MoE) Large Language Models (LLMs) deliver superior accuracy with a reduced number of active parameters, their pre-training represents a significant computationally bottleneck due to underutilized experts and limited training efficiency. This work introduces a Layer-Adaptive Expert Pruning (LAEP) algorithm designed for the pre-training stage of MoE LLMs. In contrast to previous expert pruning approaches that operate primarily in the post-training phase, the proposed algorithm enhances training efficiency by selectively pruning underutilized experts and reorganizing experts across computing devices according to token distribution statistics. Comprehensive experiments demonstrate that LAEP effectively reduces model size and substantially improves pre-training efficiency. In particular, when pre-training the Yuan3.0-1T Base model from scratch original with 1515B parameters, LAEP achieves a 48.3% improvement in training efficiency alongside a 33.3% parameter reduction, while still delivering excellent performance across multiple domains.
@@ 1 Introduction
Recent advances in Mixture-of-Experts (MoE) large language models, exemplified by models like Mixtral (Jiang et al., 2024) and DeepSeek-V3 (Liu et al., 2024a), have achieved important breakthroughs in model scale and accuracy. The MoE architecture enables a significant increase in model capacity without linearly inflating computation costs (FLOPs) by activating only a small subset of total parameters for each token, though the static parameters used to construct MoE models still take up considerable memory. During the training process, the MoE model often experiences expert load imbalance with certain experts frequently activated while some experts are rarely engaged, which can cause some experts to struggle to learn useful representations impacting the model’s overall performance and also lead to inefficient use of computational resources.
@@
  • Propose Layer-Adaptive Expert Pruning (LAEP), a novel algorithm tailored for the pre-training of MoE LLMs. Once expert loads enter the stable phase of pre-training, LAEP selectively prunes underutilized experts and rearranges the remaining experts to alleviate load imbalance across computing devices. Extensive experiments demonstrate that LAEP significantly reduces the total number of model’s parameters while improving model accuracy and pre-training efficiency.
  • Apply LAEP to pre-train a 1515B-parameter sparse MoE model, achieving a 33.3% reduction in model parameters and a 48.3% improvement in training efficiency compared with the base model. Moreover, the resulting Yuan3.0-1T pre-trained base model attains performance comparable to state-of-the-art systems on major benchmarks across diverse domains.
@@ 3.1 Expert Token distribution in Pre-training
To investigate the load characteristics of experts during the pre-training phase of MoE LLMs, we first constructed a 10B-parameter model and pre-trained it from scratch. The complete model architecture and the details of the pre-training dataset are described in Appendix A.1 and Appendix A.2, respectively.
Figure 1 illustrates the evolution of token distribution across experts at three representative layers of the 10B model over the course of pre-training. We observe that the expert token loads exhibits pronounced imbalance and follows a consistent temporal pattern across layers.
@@
Stable phase. The token distribution across experts becomes temporally stable, with per-expert token counts changing gradually over time. Despite this stability, a pronounced imbalance persists: a small subset of experts consistently processes a disproportionately large fraction of tokens, while others receive extremely few tokens and remain persistently underutilized. Without explicit load-balancing control, the disparity between the most utilized and least utilized experts can span several orders of magnitude, as shown in Figure 1. Importantly, once the stable phase is reached, the relative ordering of experts by token load becomes largely fixed. This stable ranking provides a reliable basis for identifying persistently underutilized experts.
@@ 3.2 Expert Pruning
The significant workload disparity among experts results in insufficient training for some, leading to diminished parameter efficiency and a waste of computational resources. To address this issue, we propose the LAEP method, which formulates pruning strategies adaptively at the layer level, with such strategies being determined based on the distribution of tokens across experts within each individual layer.
@@
$E^{\prime}=\text{torch.topK}\left(\sum_{j^{\prime}}E[i,j^{\prime},l],K,\text{dim}=i,\right.\\ \left.\text{largest}=\text{False}\right)$ is the topK experts with the smallest number of tokens among the experts in the $l^{th}$ layer. N denotes the number of experts in each layer before pruning. S is the total number of tokens per layer. $\alpha$ is the individual load constraint, and $\beta$ is the cumulative load constraint. Figure 2 illustrates the effects of $\alpha$ and $\beta$. After sorting experts by their token loads, any expert whose cumulative token count falls below $\beta$ of the total (Equation 3) is initially flagged corresponding to the experts located to the left of the red dashed line in Figure 2b. Among these flagged candidates, an expert is pruned if its individual token load is lower than $\alpha$ of the average load of all experts (Equation 2). The detailed implementation of expert pruning algorithm is provided in Appendix A.3.
@@ 3.3 Expert Rearrangement
To mitigate this device-level load imbalance, we propose an expert rearranging algorithm (Figure 3), which rearranges the experts across computing devices based on the expert token distribution. Starting from an imbalanced assignment of experts to computing devices, experts are first ranked according to their token loads. They are then iteratively rearranged in a greedy manner such that, at each step, the cumulative load variance across devices is minimized. By alternating the placement of high-load and low-load experts, the algorithm progressively smooths the load distribution. After rearrangement, experts are redistributed across computing devices, resulting in a significantly more balanced token load and improved utilization of computational resources. The detailed implementation of expert rearranging algorithm is provided in Algorithm 2.
@@ 3.4.1 Sensitivity Analysis of Pruning Hyper-parameters
Table 1 presents the test loss under various pruning coefficients $\alpha$ and $\beta$. When $\alpha$ is infinite, the test loss increases monotonically with $\beta$. Notably, for $\beta\leq 0.1$, the pruned model achieves a test loss of 1.658, which is lower than the base model’s 1.661. This indicates that moderate pruning can effectively enhance model performance. Motivated by this observation, we fix $\beta$ at 0.1 and vary $\alpha$ to examine its impact. While the test loss exhibits an upward trend as $\alpha$ increases, it consistently remains below the baseline for $\alpha\leq 0.4$. Under this configuration, the model achieves a 24.5% reduction in parameters compared to the base model while maintaining superior accuracy.
@@
$\beta=0.05$ 8.06 1.648
Base model pruned with $\beta=0.1$ 6.89 1.658
varied $\beta$ and $\alpha=\infty$ $\beta=0.2$ 5.10 1.679
$\beta=0.4$ 2.76 1.725
$\alpha=0.2$ 8.72 1.643
Base model pruned with $\alpha=0.2$ and 0.4 hybrid 7.89 1.650
varied $\alpha$ and $\beta=0.1$ $\alpha=0.4$ 7.38 1.653
$\alpha=0.6$ 6.51 1.661
We then examine how different auxiliary load-balancing losses influence both model accuracy and expert token distributions. Specifically, we analyze their effects on test loss (Table 1) and on the distribution of tokens across experts (Appendix A.4). We conduct experiments on the same 10B base model described in Section 3.1, comparing the baseline setting without auxiliary loss against models trained with sequence-wise auxiliary loss from DeepSeek-V3 and the auxiliary loss formulation used in Mixtral, under different weighting coefficients.
When a relatively large coefficient, such as 0.01, is applied, the auxiliary loss substantially enforces load balancing across experts, but this comes at the cost of degraded model accuracy, as reflected by a higher test loss compared to the baseline. In contrast, a smaller coefficient, such as 0.0001, yields lower test loss than the baseline, while noticeable imbalance remains: in this case, the maximum token disparity within a layer can still reach approximately 13.6×, as can be seen in Figure 4 in Appendix A.4.
@@ 3.4.2 Effect of pruning on different Attention Architectures
Table 2: Alation study on LFA with different $\alpha$.
Pruning coefficient Test Loss Test Loss w/o LFA
@@
3.4.3 Impact of LAEP on Downstream Task Performance
To validate the influence of the LAEP on the accuracy of downstream tasks, we construct a model with 20 billion parameters and conduct pre-training on a corpus of 100 billion tokens. The model’s detailed architectural specifications and the training data set are provided in Appendix A.1 and Appendix A.2.
@@ 3.4.4 Scalability of LAEP
4 Pre-training of Yuan3.0-1T Base model with LAEP
We applied LAEP to a 103-layer MoE model containing 1515 billion parameters. After pruning, the number of model layers remains unaltered, whereas the total parameter count is reduced to 1010B, representing an 33.3% reduction in parameters. This parameter optimization not only effectively alleviates the memory consumption during the pre-training phase, but also reduces the memory requirements for subsequent deployment, thereby enhancing the model’s practical applicability. The resulting model is trained from scratch on a corpus of 1.08 trillion tokens. Further details on the training data are available in Appendix A.2.
Table 5: Performance comparison of the Yuan3.0-1T Base model 1515B (activated 68.5B) under different methods. The tests are conducted on 824 AI chips, while the numerical precision during the training process is set to BF16. The coefficient for all auxiliary losses is set to 0.01, a value significantly higher than standard practice, to ensure near-ideal expert load balancing. Regarding the LAEP hyperparameters, $\alpha$ is set to 0.2 for the first and last one-sixth of the layers and 0.4 for the remaining layers, while $\beta$ is fixed at 0.1.
@@
     w/o rearrangement 103 48* 82.25
1142B DeepSeek-V3 Sequence-Wise Auxiliary Loss 103 48 84.74
@@
Table 6: Comparison on tasks between the Yuan3.0-1T Base model and other representative dense and MoE base models.
Benchmark #Shots LLaMA-3.1-405B Base DeepSeek-V3-Base Yuan3.0-1T Base
@@
Language Pile-test - 0.542 0.548 0.594
@@
As shown in Table 5, the training efficiency of the model demonstrates a significant improvement after applying LAEP. Overall computational performance increases from 62.14 TFlops per GPU to 92.6 TFlops per GPU, representing a 48.3% enhancement in training performance. Of these improvements, model pruning accounts for 32.4%, while expert rearrangement contributes 15.9%. Additionally, we compare the pre-training performance when utilizing an auxiliary loss. Notably, we employ a relatively large auxiliary loss coefficient of 0.01 to ensure near-ideal load balancing. For an analysis of the coefficient’s impact on load balancing, please refer to Figure 4 in Appendix A.4. The incorporation of the auxiliary load-balancing loss enhances training performance by 30.1%, which remains 12.7% lower than the result achieved by the proposed LAEP method. For the purpose of conducting a more comprehensive comparative study, additional experiments are designed: a 1142B model is tested under the condition where the number of experts per layer is uniformly set to 48 (while keeping the number of experts the same as the maximum number of experts that after applying LAEP on the 1515B model), and the auxiliary loss of DeepSeek-V3 is adopted in this experimental setup. The results of this control experiment demonstrate that the LAEP method still enables a 9.3% increase in training performance and an 11.6% reduction in the number of parameters, further verifying the effectiveness and superiority of the LAEP method in optimizing model performance and reducing parameter complexity.

### v3

# Yuan3.0 Ultra: A Trillion-Parameter Enterprise-Oriented MoE LLM

@@ Abstract
We introduce Yuan3.0 Ultra, an open-source Mixture-of-Experts (MoE) large language model featuring 68.8B activated parameters and 1010B total parameters, specially designed to enhance performance on enterprise scenarios tasks while maintaining competitive capabilities on general purpose tasks. We propose Layer-Adaptive Expert Pruning (LAEP) algorithm designed for the pre-training stage of MoE LLMs. In contrast to previous expert pruning approaches that operate primarily in the post-training phase, the proposed algorithm enhances training efficiency by selectively pruning underutilized experts and reorganizing experts across computing devices according to token distribution statistics. Comprehensive experiments demonstrate that LAEP effectively reduces model size and substantially improves pre-training efficiency. When pre-training Yuan3.0 Ultra from scratch original with 1515B parameters, this algorithm delivers a 49% boost in pre-training efficiency and a 33.3% reduction in total parameters, while preserving the model’s outstanding multi-domain performance. On enterprise scenario benchmarks including Docmatix, ChatRAG, SummEval and MMTab, Yuan3.0 Ultra achieves leading accuracy. The model and codes are publicly available at https://github.com/Yuan-lab-LLM/Yuan3.0-Ultra. Refer to caption Figure 1: Performance Comparison of Yuan3.0 Ultra across enterprise scenarios benchmarks.
@@ 1 Introduction
Recent advances in Mixture-of-Experts (MoE) large language models, exemplified by models like Mixtral (Jiang et al., 2024), Kimi-K2.5 (Team, 2026a), DeepSeek-V3 (Liu et al., 2024a), and Qwen3.5 (Team, 2026b), have achieved important breakthroughs in model scale and accuracy. The MoE architecture enables a significant increase in model capacity without linearly inflating computation costs (FLOPs) by activating only a small subset of total parameters for each token, though the static parameters used to construct MoE models still take up considerable memory. During the training process, the MoE model often experiences expert load imbalance with certain experts frequently activated while some experts are rarely engaged, which can cause some experts to struggle to learn useful representations impacting the model’s overall performance and also lead to inefficient use of computational resources.
@@
  • Propose Layer-Adaptive Expert Pruning (LAEP), a novel algorithm designed for the pre-training of MoE LLMs. Once expert loads enter the stable phase of pre-training, LAEP selectively prunes underutilized experts and rearranges the remaining experts to alleviate load imbalance across computing devices. Extensive experiments demonstrate that LAEP significantly reduces the total number of model’s parameters while improving model accuracy and pre-training efficiency.
  • Apply LAEP to pre-train a 1515B-parameter sparse MoE model, achieving a 33.3% reduction in model parameters and a 49% improvement in training efficiency compared with the base model. Moreover, the resulting Yuan3.0 Ultra pre-trained base model attains performance comparable to state-of-the-art systems on major benchmarks across diverse domains.
  • We refine the Reflection Inhibition Reward Mechanism (RIRM), proposed in Yuan3.0 Flash (Wu et al., 2026), and integrate it into a fast-thinking reinforcement learning paradigm for post-training. This enhanced RIRM yields a 16.33% gain in training accuracy and reduces output token length by 14.38%. As a result, Yuan3.0 Ultra achieves state-of-the-art accuracy on enterprise scenario benchmarks, and remains highly competitive in various tasks covering a wide range of domains.
@@ 3.1 Expert Token distribution in Pre-training
To investigate the load characteristics of experts during the pre-training phase of MoE LLMs, we first constructed a 20B-parameter model and pre-trained it from scratch. The complete model architecture and the details of the pre-training dataset are described in Appendix A.1 and Appendix A.2, respectively.
Figure 2 illustrates the evolution of token distribution across experts at three representative layers of the 20B model over the course of pre-training. We observe that the expert token loads exhibits pronounced imbalance and follows a consistent temporal pattern across layers.
@@
Stable phase. The token distribution across experts becomes temporally stable, with per-expert token counts changing slowly over time. Despite this stability, a pronounced imbalance persists: a small subset of experts consistently processes a disproportionately large fraction of tokens, while others receive extremely few tokens and remain persistently underutilized. Without explicit load-balancing control, the disparity between the most utilized and least utilized experts can span several orders of magnitude, as shown in Figure 2. Importantly, once the stable phase is reached, the relative ordering of experts by token load becomes largely fixed. This stable ranking provides a reliable basis for identifying persistently underutilized experts.
@@ 3.2 Expert Pruning
The significant workload disparity among experts results in insufficient training for some experts, leading to diminished parameter efficiency and a waste of computational resources. To address this issue, we propose the LAEP method, which formulates pruning strategies adaptively at the layer level, with such strategies being determined based on the distribution of tokens across experts within each individual layer.
@@
$E^{\prime}=\text{torch.topK}(\sum_{j^{\prime}}E[i,j^{\prime},l],K,\text{dim}=i,\text{largest}=\text{False})$ is the topK experts with the smallest number of tokens among the experts in the $l^{th}$ layer. N denotes the number of experts in each layer before pruning. S is the total number of tokens per layer. $\alpha$ is the individual load constraint, and $\beta$ is the cumulative load constraint. Figure 3 illustrates the effects of $\alpha$ and $\beta$ on experts pruning. After sorting experts by their token loads, any expert whose cumulative token count falls below $\beta$ of the total (Equation 3) is initially flagged as candidates to be pruned, corresponding to the experts located to the left of the red dashed line in Figure 3b. Among these flagged candidates, an expert is pruned if its individual token load is lower than $\alpha$ of the average load of all experts (Equation 2). The detailed implementation of expert pruning algorithm is provided in Appendix A.3.
@@ 3.3 Expert Rearrangement
To mitigate this device-level load imbalance, we propose an expert rearranging algorithm (Figure 4), which rearranges the experts across computing devices based on the expert token distribution. Starting from an imbalanced assignment of experts to computing devices, experts are first ranked according to their token loads. They are then iteratively rearranged in a greedy manner such that, at each step, the cumulative token variance across devices is minimized. By alternating the placement of high-load and low-load experts, the algorithm progressively smooths the load distribution. After rearrangement, experts are redistributed across computing devices, resulting in a significantly more balanced token load and improved utilization of computational resources. The detailed implementation of expert rearranging algorithm is provided in Algorithm 2.
@@ 3.4.1 Sensitivity Analysis of Pruning Hyper-parameters
Table 1 presents the test loss under various pruning coefficients $\alpha$ and $\beta$ on a 10B base model, the model architecture and details of the pre-training dataset are described in Appendix A.1 and Appendix A.2, respectively. When $\alpha$ is infinite, the test loss increases monotonically with $\beta$. Notably, for $\beta\leq 0.1$, the pruned model achieves a test loss of 1.658, which is lower than the base model’s 1.661. This indicates that moderate pruning can effectively enhance model performance. Motivated by this observation, we fix $\beta$ at 0.1 and vary $\alpha$ to examine its impact. While the test loss exhibits an upward trend as $\alpha$ increases, it consistently remains below the baseline for $\alpha\leq 0.4$. Under this configuration, the model achieves a 24.5% reduction in parameters compared to the base model while maintaining superior accuracy compared to base model.
@@
$\beta=0.05$ 8.06 1.648
Base model pruned with $\beta=0.1$ 6.89 1.658
varied $\beta$ and $\alpha=\infty$ $\beta=0.2$ 5.10 1.679
$\beta=0.4$ 2.76 1.725
$\alpha=0.2$ 8.72 1.643
Base model pruned with $\alpha=0.2$ and 0.4 hybrid 7.89 1.650
varied $\alpha$ and $\beta=0.1$ $\alpha=0.4$ 7.38 1.653
$\alpha=0.6$ 6.51 1.661
@@
Figure 5: At the stable stage of pre-training (80B tokens), the token loads of individual experts under different auxiliary loss coefficient are illustrated, with experts sorted in descending order of token count. The red dashed line denotes the average number of tokens per expert. Max/Min represents the ratio between the token count of the most heavily utilized expert and that of the least utilized expert.
We then examine how different auxiliary load-balancing losses influence both model accuracy and expert token distributions. Specifically, we analyze their effects on test loss (Table 1) and on the distribution of tokens across experts (Figure 9, Appendix A.4). We conduct experiments on the same 10B base model, comparing the baseline setting without auxiliary loss against models trained with sequence-wise auxiliary loss from DeepSeek-V3 and the auxiliary loss formulation used in Mixtral, under different weighting coefficients.
When a relatively large coefficient, such as 0.01, is applied, the auxiliary loss substantially enforces load balancing across experts as shown in Figure 5, but this comes at the cost of degraded model accuracy, as reflected by a higher test loss compared to the baseline. In contrast, a smaller coefficient, such as 0.0001, yields lower test loss than the baseline, while noticeable imbalance remains: in this case, the maximum token disparity within a layer can still reach approximately 13.6× (Figure 5).
@@ 3.4.2 Effect of pruning on different Attention Architectures
Table 2: Ablation study on LFA with different $\alpha$.
Pruning coefficient Test Loss with LFA Test Loss with Self-Attention
@@
3.4.3 Impact of LAEP on Downstream Task Performance
To validate the influence of the LAEP on the accuracy of downstream tasks, we conduct pre-training of a 20B model on a corpus of 100 billion tokens. The model’s detailed architectural specifications and the training data set are provided in Appendix A.1 and Appendix A.2.
@@ 3.4.4 Scalability of LAEP
4 Pre-training of Yuan3.0 Ultra Base model with LAEP
We applied LAEP to a 103-layer MoE model containing 1515 billion parameters. After pruning, the number of model layers remains unaltered, whereas the total parameter count is reduced to 1010B, representing an 33.3% reduction in parameters. This parameter optimization not only effectively alleviates the memory consumption during the pre-training phase, but also reduces the memory requirements for subsequent deployment, thereby enhancing the model’s practical applicability. Details on the training data are available in Appendix A.2.
Table 5: Performance comparison of the Yuan3.0 Ultra Base model 1515B (activated 68.8B) under different methods. The tests are conducted on 824 AI chips, while the numerical precision during the training process is set to BF16. The coefficient for all auxiliary losses is set to 0.01, a value significantly higher than standard practice, to ensure near-ideal expert load balancing. Regarding the LAEP hyperparameters, $\alpha$ is set to 0.2 for the first and last one-sixth of the layers and 0.4 for the remaining layers, while $\beta$ is fixed at 0.1.
@@
As shown in Table 5, the training efficiency of the model demonstrates a significant improvement after applying LAEP. Overall computational performance increases from 62.14 TFlops per GPU to 92.6 TFlops per GPU, representing a 49% enhancement in training performance. Of these improvements, model pruning accounts for 32.4%, while expert rearrangement contributes 15.9%. Additionally, we compare the pre-training performance when utilizing an auxiliary loss. Notably, we employ a relatively large auxiliary loss coefficient of 0.01 to ensure near-ideal load balancing. For an analysis of the coefficient’s impact on load balancing, please refer to Figure 5. The incorporation of the auxiliary load-balancing loss enhances training performance by 30.1%, which remains 12.7% lower than the result achieved by the proposed LAEP method.
Table 6: Comparison on tasks between the Yuan3.0 Ultra Base model and other representative dense and MoE base models.
Benchmark #Shots LLaMA-3.1-405B Base DeepSeek-V3-Base Yuan3.0 Ultra Base
@@
Language Pile-test - 0.54 0.55 0.59